OpenStudy (anonymous):

solve using substitution 1) 6x + 2y = 0 4x + y = -1 2) 3x + 2y = -4 x + 2y = -8 3) 5x - 4y = 14 x - 2y = -2 4) x + 3y = -4 4x + 5y = 5 5) x + 3y = -5 2x - y = 18 just answers please im running outa time!

OpenStudy (abb0t):

That's a lot of questions. I can help you with the first one, maybe second one but you will have to work on the others yourself.

OpenStudy (anonymous):

okay how do i start it?

OpenStudy (abb0t):

I am going to explain this as best I can. Try and follow me, if you have questions, please ask so I know you understand. Ok?

OpenStudy (anonymous):

ok thank you!

OpenStudy (abb0t):

So, the idea of substitution is to set something equal to something so that you can substitute it into another equation. Do you remember from a few chapters or pre-algebra, when you had something that looked like f(4) = x+5 and all you had to do was substitute \(4\) whenever you saw an "x"? So, f(4) when you had f(x) = x+5 \(\Rightarrow\) 4+5 = 9

OpenStudy (abb0t):

Well, this is the same thing. Except, now you are using your algebra tricks to get something in ONLY terms of "x" \(or\) "y". So that you can substitute it into equation

OpenStudy (anonymous):

yes i remember!

OpenStudy (abb0t):

Let's start by givint them numbers. (1). 6x + 2y = 0 (2). 4x + y=-1

OpenStudy (anonymous):

sooo...

OpenStudy (abb0t):

Now, rearrange (1). to get ONLY "x". So you will subtract 6x to both sides STEP 1: 6x-6x+2y=0-6x \(\Rightarrow\) 2y = -6x STEP 2: divide by 2, to get "y" by itself \(\large\frac{2}{2}y = \frac{-6}{2}x\) \(\Rightarrow\) y = -3x

OpenStudy (anonymous):

so y=-3???

OpenStudy (abb0t):

Now, you see how you have \(y = -3x\)?? Now, you can substitute your NEW (1), call it equation (1A) into equation (2).

OpenStudy (abb0t):

It is like saying f(-3x) = 4x+y=-1 except you are substituting it for \(y\) because you have \(y = -3x\)

OpenStudy (abb0t):

So you get \(4x-3x=-1\)

OpenStudy (abb0t):

\(x=-1\)

OpenStudy (anonymous):

so the answer to number 1 is x =-1?

OpenStudy (abb0t):

NOW, you have solved for "x" That means x = -1! now, plug in your found "x" into equation (1). and solve for \(y\)!! and you're done!!

OpenStudy (abb0t):

\(6(-1)+2y = 0\) solving for \(y\) you get \(-6+2y=0\) and then

OpenStudy (abb0t):

\(2y=6\) and you get y = 3! So your final answer is: \( x = -1\) and \(y = 3\) OR (-1, 3)

OpenStudy (anonymous):

sweet thank you very much!

OpenStudy (abb0t):

You can check if these are right by plugging in those values for any of the two equations for your given system! For instance, if you plug in x = -1 and y = 3 for the first equation, check it out! \(6(-1)+2(3) = 0\) and \(0=0\) is a true statement. Therefore, x = -1 and y = 3 are your answers!

OpenStudy (anonymous):

do you know any of the answers to the other ones tho?

OpenStudy (abb0t):

Yes.

OpenStudy (anonymous):

could i just have the answers now im running outa time to do this workshett!